OpenStudy (anonymous):

Match the graph with the correct system of inequalities. (systems of inequalities and graph coming)

OpenStudy (anonymous):

\[a) x ^{2} + y ^{2} \ge 9 ; y \ge x ^{2}\]

OpenStudy (anonymous):

\[b) x ^{2} + y ^{2} \ge 9 ; y \le x ^{2}\]

OpenStudy (anonymous):

\[c) x ^{2} + y ^{2} \le 9 ; y \le x ^{2}\]

OpenStudy (anonymous):

\[d) x ^{2} + y ^{2} \le 9 ; y \ge x ^{2}\]

OpenStudy (anonymous):

e) none of these